OpenStudy (anonymous):

How to find cos sin tan of any angle eg- cos 37°

OpenStudy (anonymous):

finding this will be more difficult than all of its application at your level so just remember it..

OpenStudy (cggurumanjunath):

cos(37) use a calculator !

OpenStudy (ybarrap):

This is a good place to start (see attachment above): For example, cos(30) = sqrt(3)/2 sin(30) = 1/2 tan(60) = sin(60)/cos(60) = (sqrt(3)/2) / (1/2) = sqrt(3)